# SAC: A Framework for Measuring and Inducing Personality Traits in LLMs with Dynamic Intensity Control

### v1

@@ Abstract
Large language models (LLMs) have gained significant traction across a wide range of fields in recent years. There is also a growing expectation for them to display human-like personalities during interactions. To meet this expectation, numerous studies have proposed methods for modelling LLM personalities through psychometric evaluations. However, most existing models face two major limitations: they rely on the Big Five (OCEAN) framework, which only provides coarse personality dimensions, and they lack mechanisms for controlling trait intensity. In this paper, we address this gap by extending the Machine Personality Inventory (MPI), which originally used the Big Five model, to incorporate the 16 Personality Factor (16PF) model, allowing expressive control over sixteen distinct traits. We also developed a structured framework known as Specific Attribute Control (SAC) for evaluating and dynamically inducing trait intensity in LLMs. Our method introduces adjective-based semantic anchoring to guide trait intensity expression and leverages behavioural questions across five intensity factors: Frequency, Depth, Threshold, Effort, and Willingness. Applying this methodology to three advanced LLMs - GPT-4o, Claude 3.7 Sonnet, and Gemini 2.5 Flash - we probe into the reliability of our method and test its generalizability. Through experimentation, we find that modelling intensity as a continuous spectrum yields substantially more consistent and controllable personality expression compared to binary trait toggling. Moreover, we observe that changes in target trait intensity systematically influence closely related traits in psychologically coherent directions, suggesting that LLMs internalize multi-dimensional personality structures rather than treating traits in isolation. Our work opens new pathways for controlled and nuanced human-machine interactions in domains such as healthcare, education, and interviewing processes, bringing us one step closer to truly human-like social machines.
[…]
0254
@@ 1 Introduction
Despite this, contemporary approaches to modelling personality in LLMs [7, 18, 22, 24, 27] treat it as either an emergent artifact or a binary toggle - traits are either induced or not, present or absent. This binary framing fails to reflect the psychological reality that personality traits exist along continuous dimensions [14]. Moreover, evaluations of LLM personality have been largely confined to the broad traits of the Big Five (OCEAN) [15], limiting fine-grained behavioral control and analysis.
To address these limitations, we introduce a novel framework for continuous personality induction in LLMs. Our approach leverages the 16 Personality Factor (16PF) model [9], a psychometrically grounded alternative to the Big Five, to enable expressive, fine-grained control across sixteen traits. We begin by constructing a new benchmark called PERS-16, which adapts the original MPI inventory [18] to the 16PF framework using 163 validated International Personality Item Pool (IPIP) [17] items. Then we apply P2 prompting, an established method for inducing traits in LLMs [18], to PERS-16 and show that it fails to provide reliable, fine-grained trait modulation across all traits. In response, we introduce a new framework called Specific Attribute Control (SAC), which allows traits to be induced at graded intensity levels (1-5). SAC defines personality intensity across five interpretable behavioral dimensions - Frequency, Depth, Threshold, Effort, and Willingness - and employs adjective-based semantic anchoring to stabilize trait expression. We evaluate models in both their SAC-neutral (uninduced) and SAC-induced states to test the precision and consistency of induced traits. Our experiments on three advanced LLMs - GPT-4o, Claude 3.7 Sonnet, and Gemini 2.5 Flash - show that SAC reliably shifts trait intensities in a continuous and interpretable manner. Our experiments also show that the conitnuous spectra of intensity induction in SAC is superior to binary trait induction as seen in P2 Prompting. Moreover, we observe that closely related traits consistently co-vary, revealing latent inter-trait structures that mirror human psychometric groupings.
@@ 2 Related Work and Background
Over the years, researchers have used various frameworks to assess human personality traits. The Big Five Personality Factors (OCEAN)  [28] have become a cornerstone in psychology due to their robust psychometric properties and predictive power across domains. The traits used include Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism (OCEAN) [15]. However, the Big Five framework, while broadly validated, offers a relatively coarse-grained view of personality traits [1]. In contrast, Cattell’s Sixteen Personality Factors (16PF) [9] provide a more nuanced psychometric structure, covering sixteen specific dimensions as shown in Table 1. The 16PF has seen widespread application across clinical, occupational, and research settings [2, 6, 11, 16, 26].
@@
Figure 1: Initial MPI Evaluation Prompt
Early studies on LLMs hinted that models could mimic certain personality-like attributes  [5, 13, 25, 29, 30], though these were typically small-scale demonstrations in ad hoc settings rather than standardized evaluations. A major advancement came with the Machine Personality Inventory (MPI) and P2 prompting [18], which jointly provided a psychometrically-informed framework for evaluating LLMs’ personality consistency and inducing specific traits without fine-tuning. Subsequent work [27] further validated the reliability and construct validity of inferred personality scores in LLMs.
In our work, we adopt the 16PF model, while introducing a structured intensity modulation framework, and analyzing latent trait interdependencies, which were not supported in earlier work.
@@ 3 PERS-16
To capture a more nuanced and granular understanding of personality in LLMs, we developed PERS-16 (Personality Evaluation and Rating System - 16), an extension of the original Machine Personality Inventory (MPI) adapted to Cattell’s 16 Personality Factor (16PF) model that uses 16 traits shown in Table 1. Expanding from five to sixteen traits allows PERS-16 to provide significantly more behavioural granularity, enabling a more holistic personality profile and revealing subtleties that the Big Five may overlook [1]. For instance, traits like Sensitivity and Orderliness are collapsed under broader Big Five factors (e.g., Agreeableness or Conscientiousness), but in the 16PF framework, they are independently measured, allowing for precise targeting and control. We now explain the construction of the PERS-16 dataset of psychometric questions used to evaluate state-of-the-art LLMs and compare their mean response values.
@@ 3.1 Dataset Construction and Methodology
The construction of the PERS-16 dataset was grounded in the International Personality Item Pool (IPIP) [17], a well-established, public-domain repository of psychometric items (questions). Each item corresponds to a specific personality trait in the 16PF model. In total, we used all 163 items from the IPIP pool that comprehensively cover all sixteen traits, ensuring conceptual alignment with the 16PF taxonomy. The complete list of items is provided in the supplementary materials, which are publicly accessible via Zenodo [12]. To adapt this dataset for evaluation of LLMs, we designed a consistent prompt template modeled after the original MPI framework [18]. The final prompt used can be seen in Figure 1. The {} placeholder is filled with an IPIP-derived behavioral statement, such as "enjoy bringing people together" or "know how to comfort others". Each statement is pre-labelled with its associated 16PF trait and its polarity (i.e., whether agreement indicates high or low trait presence), ensuring consistency in scoring. We implement a scoring mechanism that maps each of the 163 LLM responses to a value between 1 and 5, consistent with psychometric evaluation norms. For positively keyed items (i.e., those where agreement indicates higher trait presence), the options are scored from A = 5 to E = 1. For negatively keyed items, the scoring is reversed. The trait-wise score ${Score}_{d}$ for a given dimension (trait) $d$ is computed as the mean of responses across all items $\alpha\in IP_{d}$:
$\scriptsize Score_{d}=\frac{1}{N_{d}}\sum_{\alpha\in IP_{d}}f(\text{LLM}(\alpha,\text{template}))$ (1)
@@
where $IP_{d}$ is the item pool for trait $d$; $N_{d}$ is the number of items for trait $d$; ${LLM}(\cdot)$ returns the model’s letter-based response; $f(\cdot)$ converts this response to a numerical score (1–5), based on the polarity of the key. In addition to the mean score, we compute the variance ($\sigma^{2}$) for each trait, which serves as a proxy for internal consistency. Here, lower variance implies greater stability in how the model responds to related prompts.
Table 1: Trait means and variances for Cattell’s 16 Personality Factors across GPT-4o, Claude 3.7 Sonnet, and Gemini 2.5 Flash
Trait GPT-4o Claude Gemini 2.5 Flash SD
Mean Var. Mean Var. Mean Var.
WARMTH 4.60 0.80 4.90 0.30 4.70 0.64 0.15
INTELLECT 4.62 1.08 4.15 1.17 4.38 1.44 0.23
EMOTIONAL STABILITY 4.60 0.80 4.70 0.64 5.00 0.00 0.21
ASSERTIVENESS 3.40 0.80 3.40 1.11 3.30 1.90 0.06
GREGARIOUSNESS 2.80 0.60 3.00 0.89 2.20 1.40 0.42
DUTIFULNESS 3.40 1.20 4.20 1.08 3.40 1.69 0.46
FRIENDLINESS 3.40 1.20 3.80 0.87 3.20 1.40 0.31
SENSITIVITY 3.80 0.98 4.10 1.30 3.70 1.55 0.21
DISTRUST 2.50 0.81 1.50 0.67 2.00 1.18 0.50
IMAGINATION 3.40 1.20 3.10 0.83 2.50 1.57 0.46
RESERVE 3.60 1.56 2.40 1.28 3.10 1.81 0.60
ANXIETY 2.00 1.00 2.20 0.87 1.70 1.27 0.25
COMPLEXITY 4.60 0.80 4.90 0.30 4.60 0.66 0.17
INTROVERSION 3.20 1.08 2.30 0.78 3.60 1.36 0.67
ORDERLINESS 3.80 0.98 4.20 0.75 4.00 1.61 0.20
EMOTIONALITY 1.50 0.81 1.10 0.30 1.60 1.02 0.26
@@
Model GPT-4o Claude 3.7 Sonnet Gemini 2.5 Flash
GPT-4o 0.00 2.25 1.50
Claude 3.7 Sonnet 2.25 0.00 2.33
Gemini 2.5 Flash 1.50 2.33 0.00
3.2 Results on Neutral Model
Table 1 summarises the mean scores and variances for each trait in the baseline LLM evaluations. A clear three-tier pattern emerges: all models score highly ($\geq$ 4.5) on the “likeable/cerebral” cluster - Warmth, Intellect, Emotional Stability, and Complexity - indicating that each system is designed to appear knowledgeable, steady, and prosocial. A middle band (2 $<$ score $<$ 4.5) includes traits related to social style and work ethic - e.g., Assertiveness, Dutifulness, Orderliness - where divergence creates distinctive but balanced personalities. Finally, all models uniformly suppress Distrust, Anxiety, and Emotionality (scores $\leq$ 2), reinforcing the goal of calm, emotionally neutral assistants. Overall, the pattern suggests deliberate personality shaping rather than incidental emergence.
To contextualize LLM scores, we reference psychometric literature describing typical response patterns in human-administered 16PF assessments [9]. Most personality inventories, including the 16PF and its IPIP derivatives, tend to show mean trait scores near the midpoint of the scale (i.e., $\sim$3 on a 1-5 scale) and variances in the range of 0.8 to 1.2, though these values vary by trait and demographic sample [9]. While not using a specific norm dataset, our PERS-16 results from Table 1 exhibit similar score distributions and variances, further proving that LLMs are indeed capable of exhibiting human-like behaviour.
@@ 3.3 Cross-Model Comparison
In this subsection, we examine how different LLMs diverge in their personality behavior and explore possible causes for these differences. Our analysis focused on two aspects: (1) the Euclidean distance between trait profiles, which succinctly captures how training data and architectural choices distinguish models from one another; and (2) the standard deviation of trait scores across models, which highlights traits that are most divisive or consistently expressed - where a higher $\sigma$ indicates greater disagreement and a lower $\sigma$ suggests alignment. The Euclidean distance between two models $A$ and $B$, based on their mean personality trait scores across the 16PF traits, is defined as:
$\scriptsize d(A,B)=\sqrt{\sum_{i=1}^{16}\left(\mu_{i}^{(A)}-\mu_{i}^{(B)}\right)^{2}}$ (2)
where, $\mu_{i}^{(A)}$ is the mean score for the $i^{\text{th}}$ trait in model $A$ and $\mu_{i}^{(B)}$ is the mean score for the $i^{\text{th}}$ trait in model $B$, summed up for all sixteen traits. The results presented in Table 2 show that GPT-4o and Gemini 2.5 Flash are the most similar pair, with a Euclidean distance of 1.50. Gemini 2.5 Flash and Claude 3.7 Sonnet are the most dissimilar, with a distance of 2.33. We also observe that GPT-4o exhibits relatively low Euclidean distances to all other models, suggesting it is the most behaviorally aligned overall. Given its proximity to multiple systems, it may serve as a useful reference point for future comparisons of personality modulation across LLMs. Now, to compute the standard deviation of each trait across models, we use:
$\scriptsize\sigma_{i}=\sqrt{\frac{1}{M}\sum_{j=1}^{M}\left(\mu_{i}^{(j)}-\bar{\mu}_{i}\right)^{2}}$ (3)
where, $\mu_{i}^{(j)}$ is the mean score of the $i^{\text{th}}$ trait in the $j^{\text{th}}$ model; $\bar{\mu}_{i}=\frac{1}{M}\sum_{j=1}^{M}\mu_{i}^{(j)}$ is the mean of means for trait $i$; $M$ is the number of models (here, $M=3$); and $\sigma_{i}$ is the standard deviation for the $i^{\text{th}}$ trait across the models. The standard deviation across sixteen traits presented in Table 1 indicate that Introversion shows the highest standard deviation across models (0.67). This shows significant divergence in how LLMs represent social withdrawal and solitary tendencies. Reserve (0.60) and Distrust (0.50) also exhibit considerable variability, suggesting inconsistent modelling of social reticence and cautiousness. In contrast, Assertiveness (0.06), Warmth (0.15), and Complexity (0.17) are among the most stable traits, implying that these are represented relatively consistently across all three LLMs. The remaining ten traits show moderate standard deviation across all models, indicating that they are neither completely stable nor very inconsistent.

### v2

@@ Abstract
Large language models (LLMs) have rapidly gained traction across fields, with growing expectations for them to exhibit human-like personalities. To meet this expectation, various studies have explored personality modelling through psychometric frameworks. However, most existing approaches rely on the Big Five (OCEAN) model, which offers only broad personality dimensions and lacks mechanisms for fine-grained control of trait intensity. We address this gap by extending the Machine Personality Inventory (MPI), which originally used the Big Five model, to incorporate the 16 Personality Factor (16PF) model, allowing expressive control over sixteen distinct traits. We also developed a structured framework known as Specific Attribute Control (SAC) for evaluating and dynamically inducing trait intensity in LLMs. Our method introduces adjective-based semantic anchoring to guide trait intensity expression and leverages behavioural questions across five intensity factors: Frequency, Depth, Threshold, Effort, and Willingness. We tested this methodology on four state-of-the-art LLMs and benchmarked it against human responses, finding that modelling intensity as a continuous spectrum yields more consistent and controllable personality expression than binary trait toggling. We also find that altering a target trait’s intensity systematically shifts related traits in psychologically coherent directions, indicating that LLMs internalize multi-dimensional personality structures rather than treating traits independently. These findings open new directions for developing LLMs with adaptive, transparent, and context-sensitive personality control.
@@ 1 INTRODUCTION
Despite this, contemporary approaches to modelling personality in LLMs treat it as either an emergent artifact or a binary toggle - traits are either induced or not, present or absent. This binary framing fails to reflect the psychological reality that personality traits exist along continuous dimensions [12]. Moreover, evaluations of LLM personality have been largely confined to the broad traits of the Big Five (OCEAN) [13], limiting fine-grained behavioral control and analysis.
To address these limitations, we introduce a novel framework for continuous personality induction in LLMs. Our approach leverages the 16 Personality Factor (16PF) model [8], a psychometrically grounded alternative to the Big Five, to enable expressive, fine-grained control across sixteen traits. We begin by constructing a new benchmark called PERS-16, which adapts the original MPI inventory [16] to the 16PF framework using 163 validated International Personality Item Pool (IPIP) [15] items. Then we apply P2 prompting, an established method for inducing traits in LLMs [16], to Personality Evaluation and Rating System - 16 (PERS-16) and show that it fails to provide reliable, fine-grained trait modulation across all traits. In response, we introduce a new framework called Specific Attribute Control (SAC), which allows traits to be induced at graded intensity levels (1-5). SAC defines personality intensity across five interpretable behavioral dimensions - Frequency, Depth, Threshold, Effort, and Willingness - and employs adjective-based semantic anchoring to stabilize trait expression. We evaluate models in both their SAC-neutral (uninduced) and SAC-induced states to test the precision and consistency of induced traits. Our experiments on four advanced LLMs - GPT-4o, Claude 3.7 Sonnet, Gemini 2.5 Flash, and Mistral - show that SAC reliably shifts trait intensities in a continuous and interpretable manner. Our experiments also show that the continuous spectra of intensity induction in SAC is superior to binary trait induction as seen in P2 Prompting. Moreover, we observe that closely related traits consistently co-vary, revealing latent inter-trait structures that mirror human psychometric groupings.
@@ 2 RELATED WORK AND BACKGROUND
Over the years, researchers have used various frameworks to assess human personality traits. The Big Five Personality Factors (OCEAN)  [27] have become a cornerstone in psychology due to their robust psychometric properties and predictive power across domains. The traits used include Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism (OCEAN) [13]. Prior studies have consistently employed Likert-scale personality testing as a standard method for assessing LLM personality traits [6, 16, 22, 24, 26, 17, 20]. However, the Big Five framework, while broadly validated, offers a relatively coarse-grained view of personality traits [1]. In contrast, Cattell’s Sixteen Personality Factors (16PF) [8] provide a more nuanced psychometric structure, covering sixteen specific dimensions as shown in Table 1. The 16 Personality Factor (16PF) model has been widely employed across diverse research contexts - from comparing population group responses and examining personality-behavior relationships in clinical settings such as acute myocardial infarction, to exploring marital compatibility hypotheses and predicting psychological constructs like self-esteem using regression models derived from 16PF factor scores. [2, 5, 9, 14, 25].
@@
Figure 1: Initial MPI Evaluation Prompt.
@@ 3 PERS-16
To capture a more nuanced and granular understanding of personality in LLMs, we developed PERS-16 (Personality Evaluation and Rating System - 16), an extension of the original Machine Personality Inventory (MPI) adapted to Cattell’s 16 Personality Factor (16PF) model that uses 16 traits shown in Table 1. Expanding from five to sixteen traits allows PERS-16 to provide significantly more behavioural granularity, enabling a more holistic personality profile and revealing subtleties that the Big Five may overlook [1]. For instance, traits like Sensitivity and Orderliness are collapsed under broader Big Five factors (e.g., Agreeableness or Conscientiousness), but in the 16PF framework, they are independently measured, allowing for precise targeting and control. We now explain the construction of the PERS-16 dataset of psychometric questions used to evaluate LLMs and compare their mean response values.
Table 1: Trait means and standard deviations for Cattell’s 16 Personality Factors across GPT-4o, Claude 3.7 Sonnet, Gemini 2.5 Flash, and Mistral
Trait GPT-4o Claude 3.7 Sonnet Gemini 2.5 Flash Mistral
Mean SD Mean SD Mean SD Mean SD
WARMTH 4.60 0.80 4.90 0.30 4.70 0.64 3.60 0.66
INTELLECT 4.62 1.08 4.15 1.17 4.38 1.44 3.54 0.50
EMOTIONAL STABILITY 4.60 0.80 4.70 0.64 5.00 0.00 3.40 0.49
ASSERTIVENESS 3.40 0.80 3.40 1.11 3.30 1.90 3.40 0.66
GREGARIOUSNESS 2.80 0.60 3.00 0.89 2.20 1.40 3.30 0.64
DUTIFULNESS 3.40 1.20 4.20 1.08 3.40 1.69 3.00 0.63
FRIENDLINESS 3.40 1.20 3.80 0.87 3.20 1.40 3.10 0.83
SENSITIVITY 3.80 0.98 4.10 1.30 3.70 1.55 3.50 0.67
DISTRUST 2.50 0.81 1.50 0.67 2.00 1.18 2.80 0.75
IMAGINATION 3.40 1.20 3.10 0.83 2.50 1.57 3.50 0.67
RESERVE 3.60 1.56 2.40 1.28 3.10 1.81 2.90 0.54
ANXIETY 2.00 1.00 2.20 0.87 1.70 1.27 3.00 0.77
COMPLEXITY 4.60 0.80 4.90 0.30 4.60 0.66 3.50 0.81
INTROVERSION 3.20 1.08 2.30 0.78 3.60 1.36 3.30 0.78
ORDERLINESS 3.80 0.98 4.20 0.75 4.00 1.61 3.30 0.90
EMOTIONALITY 1.50 0.81 1.10 0.30 1.60 1.02 2.80 0.60
@@ 3.1 Dataset Construction And Methodology
The construction of the PERS-16 dataset was grounded in the International Personality Item Pool (IPIP) [15], a well-established, public-domain repository of psychometric items (questions). Each item corresponds to a specific personality trait in the 16PF model. In total, we used all 163 items from the IPIP pool that comprehensively cover all sixteen traits, ensuring conceptual alignment with the 16PF taxonomy.
The complete list of items is provided in the supplementary materials, which are publicly accessible via Zenodo [11].
@@
To adapt this dataset for evaluation of LLMs, we designed a consistent prompt template modeled after the original MPI framework [16]. The parameters used across all models were kept consistent with a Temperature of 0, top_p sampling value of 0.95, and a maximum token limit of 400. All model APIs were stateless, meaning no cache was utilized, ensuring that results were fully reproducible and unaffected by prior interactions (the order of prompts does not affect the results). The final prompt used can be seen in Figure 1. The {} placeholder is filled with an IPIP-derived behavioral statement, such as ”enjoy bringing people together” or ”know how to comfort others”. Each statement is pre-labelled with its associated 16PF trait and its polarity (i.e., whether agreement indicates high or low trait presence), ensuring consistency in scoring. We implement a scoring mechanism that maps each of the 163 LLM responses to a value between 1 and 5, consistent with psychometric evaluation norms. For positively keyed items (i.e., those where agreement indicates higher trait presence), the options are scored from A = 5 to E = 1. For negatively keyed items, the scoring is reversed. The trait-wise score ${Score}_{d}$ for a given dimension (trait) $d$ is computed as the mean of responses across all items $\alpha\in IP_{d}$:
$Score_{d}=\frac{1}{N_{d}}\sum_{\alpha\in IP_{d}}f(\text{LLM}(\alpha,\text{template}))$ (1)
@@
where $IP_{d}$ is the item pool for trait $d$; $N_{d}$ is the number of items for trait $d$; ${LLM}(\cdot)$ returns the model’s letter-based response; $f(\cdot)$ converts this response to a numerical score (1-5), based on the polarity of the key. In addition to the mean score, we compute the standard deviation ($\sigma$) for each trait, which serves as a proxy for internal consistency. Here, lower standard deviation implies greater stability in how the model responds to related prompts.
3.2 Results on Neutral Model
Table 1 summarises the mean scores and standard deviations for each trait in the baseline LLM evaluations. A clear three-tier pattern emerges: all models score highly (within the top quartile of trait scores) on the “likeable/cerebral” cluster - Warmth, Intellect, Emotional Stability, and Complexity - indicating that each system is designed to appear knowledgeable, steady, and prosocial. A middle band (2nd and 3rd quartiles of trait scores) includes traits related to social style and work ethic - e.g., Assertiveness, Dutifulness, Orderliness - where divergence creates distinctive but balanced personalities. Finally, all models uniformly suppress Distrust, Anxiety, and Emotionality (bottom quartile of trait scores), reinforcing the goal of calm, emotionally neutral assistants. Overall, the pattern suggests deliberate personality shaping rather than incidental emergence.
To contextualize LLM scores, we reference psychometric literature describing typical response patterns in human-administered 16PF assessments [8]. Most personality inventories, including the 16PF and its IPIP derivatives, tend to show mean trait scores near the midpoint of the scale (i.e., $\sim$3 on a 1-5 scale) and standard deviations in the range of 0.89 to 1.09, though these values vary by trait and demographic sample [8]. While not using a specific norm dataset, our PERS-16 results from Table 1 exhibit similar score distributions and variances, further proving that LLMs are indeed capable of exhibiting human-like behaviour.
@@ 3.2 Results on Neutral Model
Model OpenAI Claude Gemini Mistral
OpenAI 0.00 2.24 1.50 2.99
Claude 2.24 0.00 2.33 3.94
Gemini 1.50 2.33 0.00 3.54
Mistral 2.99 3.94 3.54 0.00
@@ 3.3 Cross-Model Comparison
In this subsection, we examine how different LLMs diverge in their personality behavior and explore possible causes for these differences. Our analysis focused on the Euclidean distance between trait profiles, which succinctly captures how training data and architectural choices distinguish models from one another. The Euclidean distance between two models $A$ and $B$, based on their mean personality trait scores across the 16PF traits, is defined as:
$d(A,B)=\sqrt{\sum_{i=1}^{16}\left(\mu_{i}^{(A)}-\mu_{i}^{(B)}\right)^{2}}$ (2)
where, $\mu_{i}^{(A)}$ is the mean score for the $i^{\text{th}}$ trait in model $A$ and $\mu_{i}^{(B)}$ is the mean score for the $i^{\text{th}}$ trait in model $B$, summed up for all sixteen traits. The results presented in Table 2 show that GPT-4o and Gemini 2.5 Flash are the most similar pair, with a Euclidean distance of 1.50. Other closely related models include the pairs of Claude and GPT-4o (2.24), Gemini 2.5 Flash and Claude (2.33). On the other hand, Mistral and Claude 3.7 Sonnet are the most dissimilar, with a distance of 3.94. Other dissimilar models include the pairs of Gemini 2.5 Flash and Mistral (3.54), and GPT-4o and Mistral (2.99). The lowest Euclidean distance corresponding to Mistral being 2.99 may suggest that open-source models differ from the closed-source models in their behaviour. We also observe that GPT-4o exhibits relatively low Euclidean distances to all other models, suggesting it is the most behaviorally aligned overall. Given its proximity to multiple systems, it may serve as a useful reference point for future comparisons of personality modulation across LLMs.